OpenStudy (aimmebgood):

How do you solve this: x^(3) - 5x^(2) + 6x = 0? I did it the wrong way and got x=-1 and x=6. thats way wrong, its supposed to be 2/3,3/2. how to solve this right?

OpenStudy (anonymous):

\[x^3 - 5x^2 + 6x = 0\] \[x(x^2-5x+6)=0\] \[x(x-2)(x-3)=0\] and now it should be good right?